# Improving Indigenous Language Machine Translation with Synthetic Data and Language-Specific Preprocessing

### v1

@@ 1 Introduction
Many indigenous languages face increasing risk of endangerment due to limited digital presence and scarce linguistic resources, posing significant challenges for the development of robust machine translation (MT) systems mager2023ethical; woodbury2012endangered. Languages such as Aymara and Guarani exemplify this challenge: despite being spoken by sizable communities in the Americas, their complex morphological structures and lack of large-scale parallel corpora hinder effective MT development rodriguez2022challenges. As a result, conventional MT approaches that rely on abundant supervised data often perform poorly in these settings.
Recent advances in neural machine translation (NMT), particularly multilingual pretraining and data augmentation, have opened new opportunities for low-resource MT. Community-driven efforts such as AmericasNLP have demonstrated that multilingual models combined with synthetic data generation can substantially improve translation quality for indigenous languages ebrahimi2023findings. Motivated by these findings, this work investigates the impact of synthetic parallel data augmentation on MT performance for Aymara–Spanish and Guarani–Spanish translationwoodbury2012endangered. Specifically, we augment curated datasets with synthetic sentence pairs generated using a high-capacity multilingual MT system, following the data-centric paradigm explored in projects such as MultiScript30kdriggersellis2025multiscript30kleveragingmultilingualembeddings. In addition, we include Spanish–Quechua experiments to assess whether language-specific orthographic normalization can further improve performance for morphologically rich indigenous languages.
We fine-tune the multilingual mBART model on both curated-only and synthetically augmented datasets and evaluate performance using chrF++ as the primary metric. Through controlled experiments and comparative analysis, this study aims to assess whether forward-translated synthetic data can reliably improve translation quality in low-resource indigenous language settings, contributing to ongoing efforts within the AmericasNLP initiative to support language preservation and accessibility ebrahimi2023findings.
@@ 2 Background and Related Work
Machine translation (MT) for low-resource and indigenous languages remains a persistent challenge due to limited parallel corpora, orthographic variation, under-documented grammar, and high linguistic diversity ebrahimi2023findings; mager2021findings. These constraints often render conventional supervised MT approaches ineffective, motivating research into data-efficient and transfer-based techniques.
A prominent direction in low-resource MT has been the use of data augmentation to mitigate data scarcity. Back-translation sennrich2016improving, which generates synthetic parallel data from monolingual corpora, has been shown to improve translation quality across a range of low-resource settings. In parallel, multilingual pre-trained models such as mBART50 tang2020multilingual and NLLB-200 costa2022no have enabled effective cross-lingual transfer by leveraging shared representations across many languages, achieving strong performance even with limited task-specific supervision.
Community-driven initiatives such as AmericasNLP have played a central role in advancing MT research for indigenous languages of the Americas by providing benchmark datasets, standardized evaluation protocols, and shared tasks ebrahimi2023findings; mager2021findings; ebrahimi2024findings. Results from recent AmericasNLP shared tasks indicate that the most successful systems consistently combine multilingual pre-trained models with synthetic data generation. For example, strong submissions have employed NLLB-200 with back-translation gow2023sheffield or explored multilingual transfer using models such as mBART50 and M2M-100 tonja2023enhancing. Earlier work, including IndT5 nagoudi2021indt5, further demonstrated the benefits of training models directly on indigenous language corpora.
More recently, synthetic data generation via forward translation has gained attention as a scalable alternative to back-translation for languages with scarce monolingual resources. The MultiScript30k dataset, for instance, translated Spanish captions from Multi30k into indigenous languages such as Aymara and Guarani using NLLB-200driggersellis2025multiscript30kleveragingmultilingualembeddings, illustrating the potential of high-quality synthetic parallel data to improve MT performance when paired with robust multilingual models.
@@ 3.1 Curated Datasets
We use curated parallel datasets released as part of the AmericasNLP 2023 Shared Task ebrahimi2023findings for Aymara–Spanish (aym–es), Guarani–Spanish (gn–es), and Quechua–Spanish (quy–es) translation. The Aymara dataset consists of 6,531 training and 996 development sentence pairs, while the Guarani dataset contains 26,032 training and 995 development pairs. For Quechua–Spanish, we use the largest curated dataset in the shared task, comprising 154,008 training sentence pairs and a 996-sentence development set.
@@
The curated data spans multiple domains—particularly for Guarani, which includes legal, health, and educational content—leading to substantial lexical and structural diversity. These characteristics reflect well-known challenges for machine translation of Indigenous languages, which often lack standardized orthographies and exhibit rich morphological variation mager2021findings. The curated datasets serve as the baseline for all experiments, with synthetic data added only to the training splits as described in Section 3.2.
@@ 3.2 Synthetic Data Generation
To mitigate data scarcity, we augment the curated parallel datasets with synthetic sentence pairs generated using the MultiScript30k dataset. In this pipeline, the Spanish portion of the Multi30k dataset is forward-translated into Aymara and Guarani using the NLLB-200 (3.3B) multilingual machine translation model costa2022no. This forward-translation approach produces synthetic parallel data that preserves the semantic content of the original captions while increasing linguistic diversity on the target side.
Unlike conventional back-translation approaches sennrich2016improving, our method uses Spanish as a high-resource pivot language. This design choice is consistent with recent findings from the AmericasNLP shared tasks, where synthetic data generation via multilingual pretrained models has been shown to improve translation quality for indigenous and low-resource languages ebrahimi2023findings; tonja2023enhancing; gow2023sheffield. The resulting augmented datasets increase both the volume and variability of training examples, supporting improved generalization in low-resource settings.
@@ 3.3 Preprocessing
Despite improving overall data consistency, these preprocessing steps yield limited gains in translation quality for Aymara. We hypothesize that this is due to the language’s strongly agglutinative morphology, in which grammatical information is encoded through productive suffix chains and phonemic contrasts. Generic normalization combined with subword tokenization may fragment these morphemes, motivating morpheme-aware preprocessing or segmentation strategies as future work mager2021findings.
@@ 3.4 Model Fine-tuning
We fine-tune the multilingual mBART model (facebook/mbart-large-50) using the HuggingFace Transformers library. mBART is pretrained using a denoising autoencoding objective across 50 languages, enabling effective cross-lingual transfer in low-resource translation scenarios tang2020multilingual. We use the MBart50Tokenizer with Spanish (es_XX) as the source language and the appropriate target language tags (aym_XX for Aymara and gn_XX for Guarani).
@@
For Aymara–Spanish translation, we train using a learning rate of $2\times 10^{-5}$, batch size 16, gradient accumulation of 4, and 20 epochs. For Guarani–Spanish translation, we use a learning rate of $3\times 10^{-5}$, batch size 8, gradient accumulation of 4, and 15 epochs. All models are optimized using AdamW with a cosine learning rate schedule and warm-up. Early stopping with a patience of three evaluation intervals based on validation BLEU is applied to prevent overfitting. These configurations are consistent with recent AmericasNLP submissions gow2023sheffield; tonja2023enhancing.
@@ 3.5 Evaluation Metrics
We evaluate translation quality primarily using chrF++, computed with the official AmericasNLP 2023 evaluation script. While BLEU is monitored during training for early stopping, we do not report BLEU scores in our main results due to its known limitations for morphologically rich and agglutinative languages such as Guarani, Quechua, and Aymarapopovic-2015-chrfpopovic-2017-chrf.
@@ 4.5 Comparison with AmericasNLP 2023
We compare our systems against the AmericasNLP 2023 baseline and the best per-language results reported by the University of Sheffield on the official development sets. Following prior work, we focus on chrF++ due to BLEU’s known limitations for agglutinative languagespopovic-2015-chrfpopovic-2017-chrf.
Top-performing submissions primarily relied on large multilingual models such as NLLB-200costa2022no, often paired with back-translation or other forms of synthetic data generation gow2023sheffield; tonja2023enhancing. In particular, the University of Sheffield system employed NLLB-200 with extensive back-translation and diverse parallel sources, achieving strong results across multiple language pairs gow2023sheffield. These findings highlight the effectiveness of data-centric approaches in low-resource indigenous MT.
@@ 5 Baseline Comparison
Table 3: Development-set chrF++ comparison for Aymara (aym), Guarani (gn), and Quechua (quy) on the AmericasNLP benchmarksebrahimi-etal-2023-findings. Prior results are taken from the official shared-task reports.
To contextualize our results, we compare our system against the best-performing submission reported in the AmericasNLP 2023 shared task for Guarani–Spanish translation ebrahimi2023findings. While differences in training data and augmentation strategies prevent a strictly controlled comparison, this provides a meaningful reference point within the established evaluation framework. Our results indicate that synthetic data augmentation improves performance over a curated-only baseline and yields competitive chrF++ scores relative to reported shared-task systems.Table 3 compares our dev-set chrF++ scores with the AmericasNLP 2023 baseline and the best per-language results reported by the University of Sheffield submission. Table 3 shows that synthetic data augmentation improves chrF++ across all three languages. Gains are largest for Aymara (+3.97 chrF++), followed by Guarani (+2.00), while Quechua shows smaller but consistent improvements (+0.71), reflecting its already large curated training set. Although Aymara shows the largest relative chrF++ gain, absolute performance remains low, reinforcing the need for morphology-aware modeling. While direct comparison is limited by differences in model architecture and training configurations, this provides context for the effectiveness of our synthetic data augmentation approach.
@@ 6 Discussion
The effectiveness of our approach is consistent with findings from recent AmericasNLP shared tasks, where multilingual pre-trained models combined with synthetic data generation have emerged as strong baselines for indigenous language MT. While some systems—such as the University of Sheffield submission—achieved strong performance using larger NLLB-200 models and back-translation gow2023sheffield, our results suggest that competitive improvements can also be obtained using mBART paired with forward-translated synthetic data. This highlights the flexibility of data-centric augmentation strategies across different model architectures.
Despite these encouraging results, several limitations remain. Our evaluation relies exclusively on automatic metrics, which may not fully capture cultural nuance or idiomatic correctness in indigenous languages mager2023ethical. Future work should therefore incorporate human evaluation involving proficient speakers. Additionally, further investigation into synthetic data quality, diversity, and filtering strategies—as well as alternative augmentation methods such as back-translation—may yield additional gains. Finally, incorporating explicit linguistic knowledge of Aymara and Guarani into model design or preprocessing pipelines garvin1972urbanization represents a promising direction for improving translation accuracy while ensuring ethical and community-centered deployment of MT systems mager2023ethical.

### v2

@@ 1 Introduction
Many indigenous languages face increasing risk of endangerment due to limited digital presence and scarce linguistic resources, posing significant challenges for the development of robust machine translation (MT) systems Mager et al. (2023); Woodbury (2012). Languages such as Aymara and Guarani exemplify this challenge: despite being spoken by sizable communities in the Americas, their complex morphological structures and lack of large-scale parallel corpora hinder effective MT development Rodríguez et al. (2022). As a result, conventional MT approaches that rely on abundant supervised data often perform poorly in these settings.
Recent advances in neural machine translation (NMT), particularly multilingual pretraining and data augmentation, have opened new opportunities for low-resource MT. Community-driven efforts such as AmericasNLP have demonstrated that multilingual models combined with synthetic data generation can substantially improve translation quality for indigenous languages Ebrahimi et al. (2023a). Motivated by these findings, this work investigates the impact of synthetic parallel data augmentation on MT performance for Aymara–Spanish and Guarani–Spanish translationWoodbury (2012). Specifically, we augment curated datasets with synthetic sentence pairs generated using a high-capacity multilingual MT system, following the data-centric paradigm explored in projects such as MultiScript30kDriggers-Ellis et al. (2025). In addition, we include Spanish–Quechua experiments to assess whether language-specific orthographic normalization can further improve performance for morphologically rich indigenous languages.
We fine-tune the multilingual mBART model on both curated-only and synthetically augmented datasets and evaluate performance using chrF++ as the primary metric. Through controlled experiments and comparative analysis, this study aims to assess whether forward-translated synthetic data can reliably improve translation quality in low-resource indigenous language settings, contributing to ongoing efforts within the AmericasNLP initiative to support language preservation and accessibility Ebrahimi et al. (2023a).
@@ 2 Background and Related Work
Machine translation (MT) for low-resource and indigenous languages remains a persistent challenge due to limited parallel corpora, orthographic variation, under-documented grammar, and high linguistic diversity Ebrahimi et al. (2023a); Mager et al. (2021). These constraints often render conventional supervised MT approaches ineffective, motivating research into data-efficient and transfer-based techniques.
A prominent direction in low-resource MT has been the use of data augmentation to mitigate data scarcity. Back-translation Sennrich et al. (2016), which generates synthetic parallel data from monolingual corpora, has been shown to improve translation quality across a range of low-resource settings. In parallel, multilingual pre-trained models such as mBART50 Tang and et al. (2020) and NLLB-200 Costa-Jussà et al. (2022) have enabled effective cross-lingual transfer by leveraging shared representations across many languages, achieving strong performance even with limited task-specific supervision.
Community-driven initiatives such as AmericasNLP have played a central role in advancing MT research for indigenous languages of the Americas by providing benchmark datasets, standardized evaluation protocols, and shared tasks Ebrahimi et al. (2023a); Mager et al. (2021); Ebrahimi and others (2024). Results from recent AmericasNLP shared tasks indicate that the most successful systems consistently combine multilingual pre-trained models with synthetic data generation. For example, strong submissions have employed NLLB-200 with back-translation Gow-Smith and Sánchez Villegas (2023) or explored multilingual transfer using models such as mBART50 and M2M-100 Tonja et al. (2023). Earlier work, including IndT5 Nagoudi et al. (2021), further demonstrated the benefits of training models directly on indigenous language corpora.
More recently, synthetic data generation via forward translation has gained attention as a scalable alternative to back-translation for languages with scarce monolingual resources. The MultiScript30k dataset, for instance, translated Spanish captions from Multi30k into indigenous languages such as Aymara and Guarani using NLLB-200Driggers-Ellis et al. (2025), illustrating the potential of high-quality synthetic parallel data to improve MT performance when paired with robust multilingual models.
@@ 3.1 Curated Datasets
We use curated parallel datasets released as part of the AmericasNLP 2023 Shared Task Ebrahimi et al. (2023a) for Aymara–Spanish (aym–es), Guarani–Spanish (gn–es), and Quechua–Spanish (quy–es) translation. The Aymara dataset consists of 6,531 training and 996 development sentence pairs, while the Guarani dataset contains 26,032 training and 995 development pairs. For Quechua–Spanish, we use the largest curated dataset in the shared task, comprising 154,008 training sentence pairs and a 996-sentence development set.
@@
The curated data spans multiple domains—particularly for Guarani, which includes legal, health, and educational content—leading to substantial lexical and structural diversity. These characteristics reflect well-known challenges for machine translation of Indigenous languages, which often lack standardized orthographies and exhibit rich morphological variation Mager et al. (2021). The curated datasets serve as the baseline for all experiments, with synthetic data added only to the training splits as described in Section 3.2.
@@ 3.2 Synthetic Data Generation
To mitigate data scarcity, we augment the curated parallel datasets with synthetic sentence pairs generated using the MultiScript30k dataset. In this pipeline, the Spanish portion of the Multi30k dataset is forward-translated into Aymara and Guarani using the NLLB-200 (3.3B) multilingual machine translation model Costa-Jussà et al. (2022). This forward-translation approach produces synthetic parallel data that preserves the semantic content of the original captions while increasing linguistic diversity on the target side.
Unlike conventional back-translation approaches Sennrich et al. (2016), our method uses Spanish as a high-resource pivot language. This design choice is consistent with recent findings from the AmericasNLP shared tasks, where synthetic data generation via multilingual pretrained models has been shown to improve translation quality for indigenous and low-resource languages Ebrahimi et al. (2023a); Tonja et al. (2023); Gow-Smith and Sánchez Villegas (2023). The resulting augmented datasets increase both the volume and variability of training examples, supporting improved generalization in low-resource settings.
@@ 3.3 Preprocessing
Despite improving overall data consistency, these preprocessing steps yield limited gains in translation quality for Aymara. We hypothesize that this is due to the language’s strongly agglutinative morphology, in which grammatical information is encoded through productive suffix chains and phonemic contrasts. Generic normalization combined with subword tokenization may fragment these morphemes, motivating morpheme-aware preprocessing or segmentation strategies as future work Mager et al. (2021).
@@ 3.4 Model Fine-tuning
We fine-tune the multilingual mBART model (facebook/mbart-large-50) using the HuggingFace Transformers library. mBART is pretrained using a denoising autoencoding objective across 50 languages, enabling effective cross-lingual transfer in low-resource translation scenarios Tang and et al. (2020). We use the MBart50Tokenizer with Spanish (es_XX) as the source language and the appropriate target language tags (aym_XX for Aymara and gn_XX for Guarani).
@@
For Aymara–Spanish translation, we train using a learning rate of $2\times 10^{-5}$, batch size 16, gradient accumulation of 4, and 20 epochs. For Guarani–Spanish translation, we use a learning rate of $3\times 10^{-5}$, batch size 8, gradient accumulation of 4, and 15 epochs. All models are optimized using AdamW with a cosine learning rate schedule and warm-up. Early stopping with a patience of three evaluation intervals based on validation BLEU is applied to prevent overfitting. These configurations are consistent with recent AmericasNLP submissions Gow-Smith and Sánchez Villegas (2023); Tonja et al. (2023).
@@ 3.5 Evaluation Metrics
We evaluate translation quality primarily using chrF++, computed with the official AmericasNLP 2023 evaluation script. While BLEU is monitored during training for early stopping, we do not report BLEU scores in our main results due to its known limitations for morphologically rich and agglutinative languages such as Guarani, Quechua, and AymaraPopović (2015)Popović (2017).
@@ 4.5 Comparison with AmericasNLP 2023
We compare our systems against the AmericasNLP 2023 baseline and the best per-language results reported by the University of Sheffield on the official development sets. Following prior work, we focus on chrF++ due to BLEU’s known limitations for agglutinative languagesPopović (2015)Popović (2017).
Top-performing submissions primarily relied on large multilingual models such as NLLB-200Costa-Jussà et al. (2022), often paired with back-translation or other forms of synthetic data generation Gow-Smith and Sánchez Villegas (2023); Tonja et al. (2023). In particular, the University of Sheffield system employed NLLB-200 with extensive back-translation and diverse parallel sources, achieving strong results across multiple language pairs Gow-Smith and Sánchez Villegas (2023). These findings highlight the effectiveness of data-centric approaches in low-resource indigenous MT.
@@ 5 Baseline Comparison
Table 3: Development-set chrF++ comparison for Aymara (aym), Guarani (gn), and Quechua (quy) on the AmericasNLP benchmarksEbrahimi et al. (2023b). Prior results are taken from the official shared-task reports.
To contextualize our results, we compare our system against the best-performing submission reported in the AmericasNLP 2023 shared task for Guarani–Spanish translation Ebrahimi et al. (2023a). While differences in training data and augmentation strategies prevent a strictly controlled comparison, this provides a meaningful reference point within the established evaluation framework. Our results indicate that synthetic data augmentation improves performance over a curated-only baseline and yields competitive chrF++ scores relative to reported shared-task systems.Table 3 compares our dev-set chrF++ scores with the AmericasNLP 2023 baseline and the best per-language results reported by the University of Sheffield submission. Table 3 shows that synthetic data augmentation improves chrF++ across all three languages. Gains are largest for Aymara (+3.97 chrF++), followed by Guarani (+2.00), while Quechua shows smaller but consistent improvements (+0.71), reflecting its already large curated training set. Although Aymara shows the largest relative chrF++ gain, absolute performance remains low, reinforcing the need for morphology-aware modeling. While direct comparison is limited by differences in model architecture and training configurations, this provides context for the effectiveness of our synthetic data augmentation approach.
@@ 6 Discussion
The effectiveness of our approach is consistent with findings from recent AmericasNLP shared tasks, where multilingual pre-trained models combined with synthetic data generation have emerged as strong baselines for indigenous language MT. While some systems—such as the University of Sheffield submission—achieved strong performance using larger NLLB-200 models and back-translation Gow-Smith and Sánchez Villegas (2023), our results suggest that competitive improvements can also be obtained using mBART paired with forward-translated synthetic data. This highlights the flexibility of data-centric augmentation strategies across different model architectures.
Despite these encouraging results, several limitations remain. Our evaluation relies exclusively on automatic metrics, which may not fully capture cultural nuance or idiomatic correctness in indigenous languages Mager et al. (2023). Future work should therefore incorporate human evaluation involving proficient speakers. Additionally, further investigation into synthetic data quality, diversity, and filtering strategies—as well as alternative augmentation methods such as back-translation—may yield additional gains. Finally, incorporating explicit linguistic knowledge of Aymara and Guarani into model design or preprocessing pipelines Garvin and Mathiot (1972) represents a promising direction for improving translation accuracy while ensuring ethical and community-centered deployment of MT systems Mager et al. (2023).